# Augmenting LLMs to Repair Obsolete Test Cases with Static Collector and Neural Reranker

### v1

@@ III-C Collecting TROCtx
To collect repo-level contexts, Synbciatr interacts with Language Servers via LSP [19]. Every request message conforms to the JSON-RPC-based protocol, which consists of the request type (such as ‘goto definition’) and the cursor position of the request identifier, which consists of the file path, and the indices of the line and column of the identifier within the file.
@@
For every type of TROCtx, Synbciatr first automatically analyzes the inputs and locate key identifiers (identifiers that need to find definitions and references by LSP) with their positions using the static parser tree-sitter111https://tree-sitter.github.io/tree-sitter/. The required contexts are further collected through the Language Server using the python library multilspy222https://github.com/microsoft/monitors4codegen.
III-C1 Collecting ClassCtx
@@
Based on the syntactic diff of the focal method, Synbciatr first identifies the new class types introduced in the method signature of the updated version (both parameter types and the return type will be checked). These new class types are considered as key identifiers for collecting ClassCtx. Through requesting Language Server, Synbciatr collects the definitions of these new class types as well as their parent classes. To split the collected contexts into identical chunks, Synbciatr cleans up all comments and the definition body is divided into member access operations. Only non-private declarations of fields and methods are collected. Specifically, for classes and private field accesses with Lombok annotations333https://projectlombok.org/features/(@Data, @GETTER, @SETTER), Synbciatr also retains related private field accesses.
III-C2 Collecting UsageCtx
@@
For UsageCtx, the name identifier of the updated focal method is considered as key identifier. The collecting procedures are described in Algorithm 1. On line 3, Synbciatr locates the name identifier of the updated focal in the repository. On line 4, all the usages of the updated focal method are fetched by requesting Language Server. For each usage, Synbciatr collects the usage-diff texts (diffs that contain the change of usage) in two steps, generating the diff (lines 6- 10) and gathering the required contexts (lines 11- 18). Synbciatr formats the files that contain usages before generating diff to avoid missing information. Also, since usage-diff of invocation is not sufficient for ParamSynBC and RetSynBC, Synbciatr collects additional backward and forward contexts respectively (lines 13 and 16).
@@
For the case in Fig. 1, Synbciatr finds ten usages in the repository. After filtering out usages in comments and repeated ones, four usage-diff texts are collected, one of which is demonstrated in Listing ‣ III-B.
III-C3 Collecting EnvCtx
@@
EnvCtx is constructed with additional contexts indicating environmental changes, namely env-diff texts. Synbciatr collects EnvCtx for both focal method and test method. The class file to which the method belongs and its parent classes are treated as the method’s environment. Here, the class name identifiers are the key identifiers, which are used to find out the related parent classes and collect their diffs as env-diff texts. Also, diffs are generated after cleaning comments and reformatting.
III-D Reranking TROCtx
@@
Synbciatr utilizes Neural Rerankers to filter and retain the most relevant contexts. As the test cases should be functionally consistent before and after the evolution, they have high similarity. Thus, our primary idea of rerankers is to make use of the original test to construct queries for TROCtx reranking. For the case in Fig. 1, we need to filter out unrelated member accesses and precisely retain the required method declaration “MountPOptions.getDefaultInstance()” from the collected ClassCtx of the new class MountPOptions. Given the text “MountOptions.defaults()” as the query, we can rerank the ClassCtx to get the most similar APIs with Neural Rerankers. As shown in Listing III-D, the reranking result reveals that the required API “MountPOptions.getDefaultInstance()” is ranked with the highest score under the given query.
@@
[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] java // methods are simplified as signatures public static MountPOptions getDefaultInstance(); // top1 public MountPOptions getDefaultInstanceForType(); // top2 … Reranked ClassCtx of class MountPOptions.
@@ III-D1 Constructing Queries
Based on the above ideas, the quality of the query texts decides the reranking results. However, as we have three types of TROCtxs but their formats (or granularity) are different, Synbciatr reranks each type of TROCtx with distinct queries instead of using a general one. In terms of granularity for TROCtx, ClassCtx is a fine-grained one where each context is a specific member access operation within the corresponding class (Listing ‣ III-B), while UsageCtx and EnvCtx are coarse-grained ones whose contexts directly refer to the diff texts of statements (Listings ‣ III-B and ‣ III-B). Overall, Synbciatr constructs two kinds of corresponding queries.
Constructing operation queries. We define operations as member accesses to a given class. Operation queries are used to rerank method and field declarations in ClassCtx for new class types. Based on our insight, the operations used in the original test can be reused as queries, namely operation queries, to rerank ClassCtx. For new class types in ParamSynBC and RetSynBC, Synbciatr constructs fine-grained queries by extracting operations that could potentially be accessed in the repaired test as shown in Algorithm 2. Specifically, Synbciatr constructs operation queries for new parameter and return class types respectively.
To construct operation queries for collected ClassCtx in ParamSynBC, starting from the invocation of the focal method, Synbciatr extracts operations specifically related to the obsolete parameters of the original focal method, which is deleted in the given SynBC. On line 8, Synbciatr traverses the def-use chains [22] of obsolete parameters to extract directly used operations, including methods accesses and field accesses. These operations are considered as operation queries. Besides, since the obsolete parameters may be refactored to be set by the new parameter, Synbciatr adds these operations in the form as setxxx on lines 4- 6.
To construct operation queries for collected ClassCtx in RetSynBC, starting from the invocation of the focal method, Synbciatr applies forward propagatability analysis by traversing the references of the return object of the focal method. During the analysis, related operations of the return object are extracted as the operation queries.
For the focal method in Fig. 1, only the parameter types changed. The finally extracted operation queries for the newly added parameter class type MountPOptions are as follows.
@@
Constructing statement queries. Synbciatr constructs coarse-grained queries for reranking diff texts of statements. In our settings, diff texts are collected after reformatting so that every line in the diff contains a separate statement. Since we only keep changed parts in diff texts, we should extract obsolete statements from the original test as the query, namely statement query, to rerank diff texts in UsageCtx and EnvCtx.
As recent studies have shown that LLM specializes in code summarizing and understanding tasks [23], [24], Synbciatr uses LLM to extract obsolete statements from the original test. Following existing works [25], we adopt LLMs with Few-Shot Learning and Chain-Of-Thought prompting to identify obsolete statements with the syntactic change of the focal method and original test method as inputs. LLM is first asked to summarize the syntactic differences of the focal method, and then find the obsolete test statements. Both the SynBC analysis (in natural language) and obsolete statements (stmts) are collected as coarse-grained statement queries. Specifically, to rerank the EnvCtx of the focal method, obsolete test statements are unrelated. Therefore, we use the intermediate SynBC analysis of the focal method as a coarse-grained query to search for env-diff contexts about the SynBC.
@@
SynBC analysis: “The method mount() has been updated to accept an object of type ‘MountPOptions’ instead of ‘MountOptions’ as its third parameter.”
@@
III-D2 Reranking with Queries
Synbciatr reranks each type of TROCtx based on different queries as shown in Fig. 3. For every reranking process, the top three most relevant contexts are retained by default.
To rerank ClassCtx, first, all the constructor declarations in ClassCtx are retained since they are necessary to construct instances for the class. Then, for other method and field declarations, Synbciatr transforms method declarations in ClassCtx into their signatures before reranking. According to the SynBC, ClassCtx of new parameter and return class type are reranked with and respectively.
To Rerank UsageCtx, every usage diff in UsageCtx is divided into original usage (by removing added lines in diffs) and updated usage (by removing deleted lines in diffs). Synbciatr reranks both the original and updated usages with obsolete statements as query, in which the maximum reranking score determines the relevance of the given usage diff.
To Rerank EnvCtx, EnvCtx of the focal method is directly reranked with SynBC analysis while EnvCtx of the test method is directly reranked with obsolete statements.
@@ III-E Generating Full Prompt
The full prompt consists of the unified diff of the focal method, the original test, and the related contexts (TROCtx).
Following existing works using LLM for code tasks [26], Synbciatr cleans comments in codes to avoid influencing inferring program intentions. Then, Synbciatr generates the unified diff of the focal method after formatting.
The general structure of the final prompt is demonstrated in Fig. 3. Specifically, we start by setting up the system role of LLM as an expert in Java software evolution and briefly describe the test-repair task caused by syntactic changes of the focal method. Then, we introduce the original test that needs to be repaired followed by the unified diff of the focal method. At last, constructed TROCtx is listed into different types and we ask LLM to treat these contexts as references. For each group of contexts, we also provide the basic description. For the case in Fig 1 as an example, we add a description (‘Defined in class MountPOptions’) for ClassCtx of MountPOptions.
Finally, Synbciatr requests LLM with the generated full prompt to generate the repaired test.
@@ IV-A Benchmark Datasets
The evaluation dataset, i.e., testing dataset, of Ceprot contains 520 samples extracted from 128 real-world Java projects. Based on that, we first filter out that samples with no syntactic changes in the focal method. After this step, 211 samples remain. Then, to ensure the high quality of the dataset, we manually filter noisy samples and augment new samples referring to the corresponding repository commit.
@@
We manually resolve the problems above by extracting correct co-evolution pairs with SynBC from the repository.
Augmenting new samples. To improve the diversity and generality of the dataset, we mine new samples to enrich the dataset. Specifically, we augment at most 2 samples from the real-world repository with different changing patterns for each commit in the original dataset.
To sum up, the final benchmark dataset consists of 137 samples from 54 projects. As shown in Fig. 4, the samples are diverse in the types of associated SynBC.
@@ IV-B Baselines
To assess the effectiveness of Synbciatr, we consider two baselines from existing studies and our investigation, namely Ceprot and Naivellm.
@@
Naivellm is developed based on LLM without related contexts. Compared with CodeT5, LLM is trained on huge sets of data with a much larger number of parameters. Therefore, LLM is more intelligent in general code tasks. In Naivellm, we directly use LLM to repair the test case with the focal diff (the unified diff of original and updated focal methods) and the original test method as input. Following the pre-processes in Synbciatr, comments in codes are cleaned and all the filtered codes are reformatted to a standardized style.
@@ IV-C Metrics
Textual Match. We use three specific metrics to measure how the generated code is close to the ground truth. (1) CodeBLEU. Developed from the classical machine translation evaluation metric BLEU [29], CodeBLEU [30] is widely used in evaluating code generation tasks by measuring the similarity with code semantics. Thus, we use CodeBLEU to assess the similarity between the generated method and the ground truth. (2) DiffBLEU. As only a part of statements are modified, we design the metric DiffBLEU to concentrate on the modified ones specifically. To compute it, we calculate the BLEU score of the repaired test whose unchanged statements are removed. DiffBLEU serves as a complement to CodeBLEU. (3) Accuracy (%). This metric represents the percentage of samples where the generated method is identical to the ground truth (exactly match).
@@ IV-D Implementation Settings
Synbciatr is built based on LangChain [31], which is a framework designed to simplify the procedures of developing applications powered by LLMs. With the APIs provided by LangChain, Synbciatr utilizes the SOTA LLM (GPT-4) [32] developed by openai and the widely-used open-source Neural Reranker (bge-reranker-v2-m3) [33] released by BAAI [34]. To reduce the impact of randomness, for each case in the dataset, we request LLM with the temperature as 0.1 three times and keep the best one for both Naivellm and SynBCIATR in evaluation.
@@ V Evaluation
RQ1:(Effectiveness of Synbciatr) Can Synbciatr effectively repair obsolete tests caused by SynBC?
RQ2:(Effectiveness of TROCtx) To what extent does the TROCtx contribute to correctly repairing the test?
RQ3:(Failure Analysis) Under which cases does Synbciatr fail to repair?
RQ4:(Efficiency) What is the efficiency of Synbciatr?
V-A RQ1: Effectiveness of Synbciatr
@@ V-A1 Basic Evaluation on Effectiveness
To evaluate the performance of Synbciatr compared with baselines, we adopt different approaches to repair obsolete test cases in the benchmark dataset.
First, we conduct syntax validation on the generated test codes. As shown in Tab. I, the Syntax Pass Rate (SPR) of Ceprot is 47.8%, while LLM-based approaches all pass the validation (100%), which means that LLM demonstrates a higher proficiency in producing syntactically accurate code than CodeT5. Then we use the three textual metrics to measure the average similarity between the generated test code and the ground truth. The column Textual Match in Tab. I demonstrates that our approach outperforms baselines in terms of all three metrics. Specifically, LLM-based approaches have average improvements of 12.4% and 72.4% in terms of CodeBLEU and DiffBLEU compared with Ceprot, which indicates that LLM is more capable of understanding the semantics of code and generating repaired tests. With the highest scores on CodeBLEU and DiffBLEU, Synbciatr is also able to accurately repair 32.4% of the test cases, which also achieves varying degrees of improvement over CEPROT (4.4%) and NAIVELLM (28.7%).

### v2

@@ III-C Collecting TROCtxs
To collect repo-level contexts, Synter interacts with language servers via LSP [15]. Every request message conforms to the JSON-RPC-based protocol, which consists of the request type (such as ‘goto definition’) and the cursor position of the request identifier, including the file path and the indices of the line and column for the identifier within the file.
@@
For every type of TROCtx, Synter first automatically analyzes the inputs and locate key identifiers (identifiers that need to find definitions and references by LSP) with their positions using the static parser tree-sitter111https://tree-sitter.github.io/tree-sitter/. The required contexts are further collected through the language server using the python library multilspy222https://github.com/microsoft/monitors4codegen.
III-C1 Collecting ClassCtxs
@@
Based on the syntactic diff of the focal method, Synter first identifies the new class types introduced in the method signature of the updated version (both parameter types and the return type will be checked). These new class types are considered as key identifiers for collecting ClassCtxs. Through requesting the language server, Synter collects the definitions of these new class types as well as their parent classes. To split the collected contexts into identical chunks, Synter cleans up all comments and the definition body is divided into member access operations. Only non-private declarations of fields and methods are collected. Specifically, for classes and private field accesses with Lombok annotations333https://projectlombok.org/features/(@Data, @GETTER, @SETTER), Synter also retains their related private field accesses.
III-C2 Collecting UsageCtxs
@@
For UsageCtxs, the name identifier of the updated focal method is considered as the key identifier. The collecting procedures are described in Algorithm 1. On line 3, Synter locates the name identifier of the updated focal in the repository. On line 4, all the usages of the updated focal method are fetched by requesting the language server. For each usage, Synter collects the usage-diff texts (diffs that contain the change of usage) in two steps, generating the diff (lines 6- 10) and gathering the required contexts (lines 11- 18). Synter formats the files that contain usages before generating diff to avoid missing information. Also, since usage-diff of invocation is not sufficient for ParamSynBCs and RetSynBCs, Synter collects additional backward and forward surrounding contexts respectively (lines 13 and 16).
@@
For the case in Fig. 1, Synter finds ten usages in the repository. After filtering out usages in comments and repeated ones, four usage-diff texts are collected, one of which is demonstrated in Listing ‣ III-B.
III-C3 Collecting EnvCtxs
@@
EnvCtxs represent additional contexts indicating environmental changes, namely env-diff texts. Synter collects EnvCtxs for the focal and test method respectively. Specifically, the class file to which the method belongs and its parent classes are treated as the method’s environment. Here, the class name identifiers are the key identifiers, which are used to find out the related parent classes and collect their diffs as env-diff texts. Also, diffs are generated after cleaning comments and reformatting.
III-D Reranking TROCtxs
@@
Synter utilizes neural rerankers to filter and retain the most relevant contexts. As the test cases should be functionally consistent before and after the evolution, they have high similarity. Thus, our primary idea of rerankers is to make use of the original test to construct queries for TROCtx reranking. For the case in Fig. 1, from the collected ClassCtxs of the new class MountPOptions, we hope to filter out unrelated member accesses and precisely retain the required method declaration “MountPOptions.getDefaultInstance()”. Given the text ‘MountOptions.defaults()” as a query, we can rerank the ClassCtxs to get the most similar APIs with neural rerankers. As shown in Listing III-D, the reranking result reveals that the required API “MountPOptions.getDefaultInstance()” is ranked with the highest score under the given query.
@@
[!ht] {minted} [ frame=single, framesep=2mm, baselinestretch=1, fontsize=, breaklines=true ] java // methods are simplified as signatures public static MountPOptions getDefaultInstance(); // top1 public MountPOptions getDefaultInstanceForType(); // top2 … Reranked ClassCtxs of class MountPOptions.
@@ III-D1 Constructing Queries
Based on the above ideas, the quality of the query texts decides the reranking results. However, as we have three types of TROCtxs but their formats (or granularity) are different, Synter reranks each type of TROCtxs with distinct queries instead of using a general one. In terms of granularity for TROCtx, ClassCtx is a fine-grained one where each context is a specific member access operation within the corresponding class (Listing ‣ III-B), while UsageCtx and EnvCtx are coarse-grained ones whose contexts directly refer to the diff texts of statements (Listings ‣ III-B and ‣ III-B). Overall, Synter constructs two kinds of corresponding queries, operation queries and statement queries.
Constructing operation queries. We define operation as member access to a given class. Operation queries are used to rerank method and field declarations in ClassCtxs for new class types. Based on our insight, the operations used in the original test can be reused as queries, namely operation queries, to rerank ClassCtxs. For new class types in ParamSynBCs and RetSynBCs, Synter constructs fine-grained queries by extracting operations that could potentially be accessed in the repaired test as shown in Algorithm 2. Specifically, Synter constructs operation queries for new parameter and return class types respectively.
To construct operation queries for collected ClassCtxs in ParamSynBCs, starting from the invocation of the focal method in the original test, Synter extracts backward operations related to the obsolete parameters, which are modified during the given SynBC. Since the obsolete parameters may be refactored to be set by new parameters, Synter first adds additional operation queries in the form as set_xxx on lines 4- 6. Then, on line 8, Synter traverses the def-use chains [23] of the obsolete parameters to extract directly used operations, including method accesses and field accesses. These operations are also collected as operation queries.
To construct operation queries for collected ClassCtxs in RetSynBCs, starting from the invocation of the focal method in the original test, Synter applies forward propagatability analysis by traversing the references of the return object for the focal method. During the analysis, related operations of the return object are extracted as the operation queries.
For the focal method in Fig. 1, only the parameter types changed. Therefore, Synter only extracts operation queries for the newly added parameter class type MountPOptions, which is shown as follows.
@@
Constructing statement queries. Synter constructs coarse-grained queries for reranking diff texts of statements. In our settings, diff texts are collected after reformatting so that every line in the diff contains a separate statement. Since we only keep changed parts in diff texts, we should extract obsolete statements from the original test as the query, namely statement query, to rerank diff texts in UsageCtxs and EnvCtxs.
As recent studies have shown that LLM specializes in code summarizing and understanding tasks [24], [25], Synter uses an LLM to extract obsolete statements from the original test. Following existing works [26, 27], we adopt LLMs with Few-Shots Learning and Chain-Of-Thought prompting to identify obsolete statements with the syntactic change of the focal method and original test method as input. The LLM is first asked to summarize the syntactic differences of the focal method, and then find the obsolete test statements. Both the SynBC behavior description (in natural language) and obsolete statements (stmts) are collected as coarse-grained statement queries. Specifically, the SynBC behavior description is used as coarse queries for reranking the EnvCtxs of the focal method, while the EnvCtxs of the test method are reranked with the obsolete statements.
@@
SynBC behavior description: “The method mount() has been updated to accept an object of type ‘MountPOptions’ instead of ‘MountOptions’ as its third parameter.”
@@
III-D2 Reranking TROCtxs with Queries
Synter reranks each type of TROCtxs with different queries as shown in Fig. 3. During a single reranking process, the top three most relevant contexts are retained by default.
To rerank ClassCtxs, first, all the constructor declarations in ClassCtxs are retained since they are necessary to construct instances for the class. Then, for other method and field declarations, Synter transforms method declarations in ClassCtxs into their signatures before reranking. According to the SynBC, ClassCtxs of new parameter and return class types are reranked with and respectively.
To rerank UsageCtxs, every usage-diff text in the UsageCtxs is divided into original usage (by removing added lines in diffs) and updated usage (by removing deleted lines in diffs). Synter reranks both the original and updated usages with obsolete statements as query, in which the maximum reranking score determines the relevance of the given usage-diff text.
To rerank EnvCtxs, EnvCtxs of the focal method are directly reranked with the SynBC behavior description, while EnvCtxs of the test method are reranked with the obsolete statements.
@@ III-E Generating Full Prompt
The full prompt consists of the unified diff of the focal method, the original test, and the related contexts (TROCtxs).
Following existing works using LLM for code tasks [28, 27], Synter cleans comments in codes to avoid influencing inferring program intentions. Then, Synter generates the unified diff of the focal method after formatting.
The general structure of the final prompt is demonstrated in Fig. 3. Specifically, we start by setting up the system role of LLM as an expert in Java software evolution and briefly describe the test-repair task caused by syntactic changes of the focal method. Then, we introduce the original test that needs to be repaired followed by the unified diff of the focal method. At last, we categorize TROCtxs according to their types and ask LLM to treat these contexts as references. For each group of contexts, we also provide the basic description. For the case in Fig 1 as an example, we automatically generate a description (‘Defined in class MountPOptions’) for ClassCtxs of MountPOptions.
Finally, Synter requests the LLM with the generated full prompt to generate the repaired test.
@@ IV-A Benchmark Datasets
The evaluation dataset, i.e., testing dataset, of Ceprot contains 520 samples extracted from 128 real-world Java projects. Based on that, we first filter out the samples without syntactic changes in the focal method. After this step, 211 samples remain. Then, to ensure the high quality of the dataset, we manually filter noisy samples and augment new samples referring to the corresponding repository commit.
@@
We manually resolve the problems above by extracting correct co-evolution pairs with SynBCs from the repository.
Augmenting new samples. To improve the diversity and generality of the dataset, we mine new samples to enrich the dataset. Specifically, we augment at most two samples from the real-world repository with different changing patterns for each commit in the original dataset.
To sum up, the final benchmark dataset consists of 137 samples from 54 projects. As shown in Fig. 4, the samples are diverse in the types of associated SynBCs.
@@ IV-B Baselines
To assess the effectiveness of Synter, we consider two baselines from existing studies and our investigation, namely Ceprot and Naivellm.
@@
Naivellm is developed based on LLM without related contexts. Compared with CodeT5, LLM is trained on huge sets of data with a much larger number of parameters. Therefore, LLM is more intelligent in general code tasks. In Naivellm, we directly use LLM to repair the test case with the focal diff (the unified diff of original and updated focal methods) and the original test method as input. Following the pre-processes in Synter, comments in codes are cleaned and all the filtered codes are reformatted to a standardized style.
@@ IV-C Metrics
Textual Match. We use three specific metrics to measure how the generated code is close to the ground truth. (1) CodeBLEU. Developed from the classical machine translation evaluation metric BLEU [31], CodeBLEU [32] is widely used in evaluating code generation tasks by measuring the similarity with code semantics. Thus, we use CodeBLEU to assess the similarity between the generated method and the ground truth. (2) DiffBLEU. As only a part of statements are modified, we design the metric DiffBLEU to concentrate on the modified ones specifically. To compute it, we calculate the BLEU score of the repaired test whose unchanged statements are removed. DiffBLEU serves as a complement to CodeBLEU. (3) Accuracy (%). This metric represents the percentage of samples where the generated method is identical to the ground truth (exactly match textually).
@@ IV-D Implementation Settings
Synter is built based on LangChain [33], which is a framework designed to simplify the procedures of developing applications powered by LLMs. With the APIs provided by LangChain, Synter utilizes the SOTA LLM (GPT-4) [34] developed by OpenAI and the widely-used open-source neural reranker (bge-reranker-v2-m3) [35] released by BAAI [36]. To reduce the impact of randomness, for each case in the dataset, we request LLM with the temperature as 0.1 three times and keep the best one for both Naivellm and Synter in evaluation.
@@ V Evaluation
RQ1:(Effectiveness of Synter) Can Synter effectively repair obsolete tests caused by SynBCs?
RQ2:(Effectiveness of TROCtx) To what extent do the TROCtxs contribute to correctly repairing the test?
RQ3:(Failure Analysis) Under which cases does Synter fail to repair?
RQ4:(Efficiency) What is the efficiency of Synter?
V-A RQ1: Effectiveness of Synter
@@ V-A1 Basic Evaluation on Effectiveness
To evaluate the performance of Synter compared with baselines, we adopt different approaches to repair obsolete test cases in the benchmark dataset.
First, we conduct syntax validation on the generated test codes. As shown in Tab. I, the Syntax Pass Rate (SPR) of Ceprot is 47.8%, while LLM-based approaches all pass the validation (100%), which means that LLM demonstrates a higher proficiency in producing syntactically accurate code than CodeT5. Then we use the three textual metrics to measure the average similarity between the generated test code and the ground truth. The column Textual Match in Tab. I demonstrates that our approach outperforms baselines in terms of all three metrics. Specifically, LLM-based approaches have average improvements of 12.4% and 72.4% in terms of CodeBLEU and DiffBLEU compared with Ceprot, which indicates that LLM is more capable of understanding the semantics of code and generating repaired tests. With the highest scores on CodeBLEU and DiffBLEU, Synter is also able to accurately repair 32.4% of the test cases, which also achieves varying degrees of improvement over CEPROT (4.4%) and NAIVELLM (28.7%).